# Minimax rates of convergence for nonparametric regression under adversarial attacks

### v1

@@ Abstract
Recent research shows the susceptibility of machine learning models to adversarial attacks, wherein minor but maliciously chosen perturbations of the input can significantly degrade model performance. In this paper, we theoretically analyse the limits of robustness against such adversarial attacks in a nonparametric regression setting, by examining the minimax rates of convergence in an adversarial sup-norm. Our work reveals that the minimax rate under adversarial attacks in the input is the same as sum of two terms: one represents the minimax rate in the standard setting without adversarial attacks, and the other reflects the maximum deviation of the true regression function value within the target function class when subjected to the input perturbations. The optimal rates under the adversarial setup can be achieved by a plug-in procedure constructed from a minimax optimal estimator in the corresponding standard setting. Two specific examples are given to illustrate the established minimax results.
@@ 1 Introduction
This paper determines the sup-norm rate of convergence in a nonparametric regression setup with additive perturbations, in which the attacker can add arbitrary perturbations in a set to the input, thereby degrading the performance of the trained estimator. We establish that under general class of regression functions and adversarial perturbation sets, the minimax risk converges at the order of the rate in the standard setup without adversaries, plus the maximum deviation of true function values within the target regression function class. The optimal rate can be achieved by a plug-in procedure constructed from a minimax optimal estimator in the standard setting. We provide minimax results for two specific examples of function classes, including isotropic Hölder class and anisotropic Hölder class, and investigate the effects of $\ell_{p}$-attacks ($0<p\leq\infty$) and sparse attacks under these two function classes, respectively.
@@ 1.2 Outline
The rest of this paper is organized as follows. Section 2 gives a setup for the nonparemetric regression problem and the definition of adversarial loss/risk. In Section 3, we state upper and lower bounds on the minimax risks under the adversarial attack. Two specific examples are discussed in Section 4. The full proof can be found in the Appendix.
@@ 2 Problem setup
In this paper, we consider the estimation of the regression function in the presence of an adversary. Specifically, when assessing the performance of the estimator $\widehat{f}$, the adversary can add any perturbation $\delta\in\Delta_{n}$ to the input $x$, where $\Delta_{n}\in\mathbb{R}^{d}$ is a closed set containing $\delta=0$, and $\Delta_{n}$ may depend on the sample size $n$. A representative example of $\Delta_{n}$ is the $\ell_{p}$-ball ($p>0$) $B_{p}^{q_{n}}=\{z:\|z\|_{p}\leq q_{n}\}$ centering at origin with radius $q_{n}>0$. In the adversarial setting, the sup-norm loss of estimation is defined as
$L_{\Delta_{n}}(f,\widehat{f})=\sup_{x\in\mathcal{X}}\sup_{\delta\in\Delta_{n}}\left|f(x)-\widehat{f}(x+\delta)\right|,$ (2.2)
@@
Throughout this paper, let $\mathbb{N}_{0}$ denote the set of non-negative integers. For any $a\in\mathbb{R}^{d}$ and $B\subseteq\mathbb{R}^{d}$, we use the Minkowski sum notations $a+B\triangleq\{a+b:b\in B\}$ and $a-B\triangleq\{a-b:b\in B\}$. For any positive sequences $a_{n}$ and $b_{n}$, we denote $a_{n}=O(b_{n})$ and $a_{n}\lesssim b_{n}$ if there exist $C>0$ and $N>0$ such that $n\geq N$ implies $a_{n}\leq Cb_{n}$. If $a_{n}=O(b_{n})$ and $b_{n}=O(a_{n})$, then we write $a_{n}\asymp b_{n}$.
@@ 3.1 Ideal adversarial loss
Define $\mathcal{X}^{\prime}\triangleq\cup_{x\in\mathcal{X}}(x+\Delta_{n})$. Then for any estimator $\widehat{f}$, we have
$L_{\Delta_{n}}(f,\widehat{f})=\sup_{x\in\mathcal{X}}\sup_{x^{\prime}\in x+\Delta_{n}}\left|f(x)-\widehat{f}(x^{\prime})\right|=\sup_{x^{\prime}\in\mathcal{X}^{\prime}}\sup_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}\left|f(x)-\widehat{f}(x^{\prime})\right|.$ (3.1)
@@
$L_{\Delta_{n}}^{*}(f)\triangleq\inf_{f^{\prime}}L_{\Delta_{n}}(f,f^{\prime})=\frac{1}{2}\sup_{x^{\prime}\in\mathcal{X}^{\prime}}\left[\sup_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)-\inf_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)\right],$ (3.2)
@@
$f^{*}(x)=\frac{1}{2}\left[\sup_{x^{\prime}\in(x-\Delta_{n})\cap\mathcal{X}}f(x^{\prime})+\inf_{x^{\prime}\in(x-\Delta_{n})\cap\mathcal{X}}f(x^{\prime})\right],\quad x\in\mathcal{X}^{\prime}.$ (3.3)
@@ 3.2 Minimax rates of convergence
In this subsection, our aim is to establish the minimax rates of convergence for the sup-norm risk under the adversarial attacks. We propose a plug-in procedure to achieve the minimax optimal rates, which is derived from a minimax optimal estimator in the corresponding standard setting.
@@
Subsequently, we formulate a plug-in estimator:
$\widehat{f}_{\mathrm{PI}}(x)=\frac{1}{2}\left[\sup_{x^{\prime}\in(x-\Delta_{n})\cap\mathcal{X}}\widetilde{f}(x^{\prime})+\inf_{x^{\prime}\in(x-\Delta_{n})\cap\mathcal{X}}\widetilde{f}(x^{\prime})\right],\quad x\in\mathcal{X}^{\prime}.$ (3.4)
The performance of the plug-in estimator $\widehat{f}_{\mathrm{PI}}(x)$ clearly depends on the construction of $\widetilde{f}$. The following theorem first provides an upper bound for the adversarial risk of $\widehat{f}_{\mathrm{PI}}(x)$ considering a general $\widetilde{f}$. Additionally, Theorem 2 establishes minimax upper bounds when specific choices of $\widetilde{f}$ are adopted.
@@
$R_{\Delta_{n}}(f,\widehat{f}_{\mathrm{PI}})\leq R(f,\widetilde{f})+3L_{\Delta_{n}}^{*}(f),$ (3.5)
@@
The following lower bound results show that the plug-in estimator based on $\widetilde{f}$ with (3.6) is in fact minimax rate optimal.
@@
Theorems 23 together establish the minimax rates of convergence for nonparametric regression under the adversarial attacks,
@@
If $\widetilde{f}$ is minimax optimal in the sense that $\sup_{f\in\mathcal{F}}R(f,\widetilde{f})\asymp\inf_{\widehat{f}}\sup_{f\in\mathcal{F}}R(f,\widehat{f})$ under the standard setting, then the plug-in estimator $\widehat{f}_{\mathrm{PI}}$ based on $\widetilde{f}$ is minimax optimal in terms of the adverarial risk. To the best our knowledge, (3.10) is the first minimax result for the general regression setting. Our bounds are modular and general. They can be applied to many models. Applying it to new models is easy: it requires simply computing the sup-norm convergence and the ideal adversarial loss in the target function class.
@@ 4 Applications
The marginal distribution $\mathbb{P}_{X}$ has a continuous density function that is lower bounded away from 0 on $\mathcal{X}$.
Assumption 1 is a commonly used assumption in the nonparametric regression problems with random design; for example, see Condition 3’ of Stone, (1982) and Definition 2.2 of Audibert and Tsybakov, (2007). In addition, we further assume that the random error term is distributed according to a centered Gaussian distribution, which is the scenario where the known minimax theory in sup-norm can apply (see, e.g., Stone,, 1982; Bertin, 2004b, ; Gaïffas,, 2007).
@@ 4.1 Isotropic Hölder class
$\inf_{\widehat{f}}\sup_{f\in\mathcal{F}_{1}(\beta,L)}R_{\Delta_{n}}(f,\widehat{f})\asymp\left(\frac{\log n}{n}\right)^{\frac{\beta}{2\beta+d}}+r_{n}^{1\wedge\beta}.$ (4.3)
@@
The determination of the rate of $\sup_{f\in\mathcal{F}_{1}(\beta,L)}L_{\Delta_{n}}^{*}(f)$ is provided in Section A.2.1 of the Appendix.
@@
The quantity $r_{n}$ in (4.2) measures the length of the longest line segment contained in the set $\Delta_{n}$, and it may depend on the sample size $n$. The condition imposed on $\Delta_{n}$ is quite mild, which is satisfied by the $\ell_{p}$-ball: $B_{p}^{q_{n}}\triangleq\{\delta\in\mathbb{R}^{d}:\|\delta\|_{p}\leq q_{n}\}$, $0<p\leq\infty$, and the $\ell_{p}$-ball with the $\ell_{0}$-constraint: $B_{p}^{q_{n}}\cap\{\delta:\|\delta\|_{0}\leq s_{n}\}$. Note that there is an extensive body of prior work studying adversarial machine learning based on $\ell_{0}$ (Delgosha et al.,, 2024), $\ell_{2}$ (Bhattacharjee and Chaudhuri,, 2020; Bhattacharjee et al.,, 2021), and $\ell_{\infty}$ attacks (Athalye et al.,, 2018; Marzi et al.,, 2018). However, these analyses focus on the specific attacks and lack general applicability. In contrast, the result in Example 1 sheds theoretical insight into the adversarial robustness under the general $\ell_{p}$-attacks with $0<p\leq\infty$. Specifically, when $\Delta_{n}=B_{p}^{q_{n}}$, we have $r_{n}=q_{n}$, and thus the minimax adversarial risk is given by
@@
$\inf_{\widehat{f}}\sup_{f\in\mathcal{F}_{1}(\beta,L)}R_{B_{p}^{q_{n}}}(f,\widehat{f})\asymp\left(\frac{\log n}{n}\right)^{\frac{\beta}{2\beta+d}}+{q_{n}}^{1\wedge\beta},$ (4.5)
@@
which can be reached by the plug-in estimator (3.4) with $\widetilde{f}$ constructed by a suitably designed local polynomial estimator (see, e.g., Stone, (1982), Gaïffas, (2007), and Tsybakov, (2008)).
@@
The equation (4.5) shows that when $\beta<1$ and $q_{n}\lesssim(\log n/n)^{1/(2\beta+d)}$, the minimax rate in the adversarial sup-norm remains unchanged to the standard minimax rate (4.4). However, as the magnitude of perturbation increases, e.g., $q_{n}\gtrsim(\log n/n)^{1/(2\beta+d)}$, the minimax risk has the order $q_{n}^{\beta}$. When $\beta\geq 1$ and the functions in $\mathcal{F}_{1}(\beta,L)$ become smoother, the radius for this phase transition phenomenon is $(\log n/n)^{\beta/(2\beta+d)}$. It is also worthy noting that the parameter $p$, which controls the sparsity structure of the perturbation set $B_{p}^{q_{n}}$, does not affect the adversarial minimax rates in this example.

### v2

@@ Abstract
Recent research shows the susceptibility of machine learning models to adversarial attacks, wherein minor but maliciously chosen perturbations of the input can significantly degrade model performance. In this paper, we theoretically analyse the limits of robustness against such adversarial attacks in a nonparametric regression setting, by examining the minimax rates of convergence in an adversarial sup-norm. Our work reveals that the minimax rate under adversarial attacks in the input is the same as sum of two terms: one represents the minimax rate in the standard setting without adversarial attacks, and the other reflects the maximum deviation of the true regression function value within the target function class when subjected to the input perturbations. The optimal rates under the adversarial setup can be achieved by an adversarial plug-in procedure constructed from a minimax optimal estimator in the corresponding standard setting. Two specific examples are given to illustrate the established minimax results.
@@ 1 Introduction
This paper determines the sup-norm rate of convergence in a nonparametric regression setup with additive perturbations, in which the attacker can add arbitrary perturbations in a set to the input, thereby degrading the performance of the trained estimator. We establish that under general class of regression functions and adversarial perturbation sets, the minimax risk converges at the order of the rate in the standard setup without adversaries, plus the maximum deviation of true function values within the target regression function class. The optimal rate can be achieved by an adversarial plug-in procedure constructed from a minimax optimal estimator in the standard setting. We provide minimax results for two specific examples of function classes, including isotropic Hölder class and anisotropic Hölder class, and investigate the effects of $\ell_{p}$-attacks ($0<p\leq\infty$) and sparse attacks under these two function classes, respectively.
@@ 1.2 Outline
The rest of this paper is organized as follows. Section 2 gives a setup for the nonparametric regression problem and the definition of adversarial loss/risk. In Section 3, we state upper and lower bounds on the minimax risks under the adversarial attack. Two specific examples are discussed in Section 4. Section 5 presents numerical simulation results. The proofs of the main theorems and examples are provided in the Appendix.
@@ 2 Problem setup
In this paper, we consider the estimation of the regression function in the presence of an adversary. Specifically, when assessing the performance of the estimator $\widehat{f}$, the adversary can add any perturbation $\delta\in\Delta_{n}$ to the input $x$, where $\Delta_{n}\in\mathbb{R}^{d}$ is a closed set containing $\delta=0$, and $\Delta_{n}$ may depend on the sample size $n$. A representative example of $\Delta_{n}$ is the $\ell_{p}$-ball $B_{p}^{q_{n}}=\{z:\|z\|_{p}\leq q_{n}\}$ centering at origin with radius $q_{n}>0$ and $p>0$. In the adversarial setting, the sup-norm loss of estimation is defined as
$L_{\Delta_{n}}(f,\widehat{f})=\sup_{x\in\mathcal{X}}\sup_{\begin{subarray}{c}\delta\in\Delta_{n}\\ x+\delta\in\mathcal{X}\end{subarray}}\left|f(x)-\widehat{f}(x+\delta)\right|,$ (2.2)
@@
Throughout this paper, let $\mathbb{N}_{0}$ denote the set of non-negative integers. For any $a\in\mathbb{R}^{d}$ and $B\subseteq\mathbb{R}^{d}$, we use the Minkowski sum notations $a+B\triangleq\{a+b:b\in B\}$ and $a-B\triangleq\{a-b:b\in B\}$. For any positive sequences $a_{n}$ and $b_{n}$, we denote $a_{n}=O(b_{n})$ and $a_{n}\lesssim b_{n}$ if there exist $C>0$ and $N>0$ such that $n\geq N$ implies $a_{n}\leq Cb_{n}$. If $a_{n}=O(b_{n})$ and $b_{n}=O(a_{n})$, then we write $a_{n}\asymp b_{n}$. For $1\leq p<\infty$, we use $\|\delta\|_{p}$ to denote the $\ell_{p}$-norm $(\sum_{j=1}^{d}|\delta|_{j}^{p})^{1/p}$ of a vector $\delta\in\mathbb{R}^{d}$. We use $\|\delta\|_{\infty}$ to denote the sup-norm $\sup_{1\leq j\leq d}|\delta_{j}|$. For brevity, we write $\|\delta\|$ to represent the $\ell_{2}$-norm.
@@ 3.1 Ideal adversarial loss
For any estimator $\widehat{f}$, we have
$L_{\Delta_{n}}(f,\widehat{f})=\sup_{x\in\mathcal{X}}\sup_{x^{\prime}\in(x+\Delta_{n})\cap\mathcal{X}}\left|f(x)-\widehat{f}(x^{\prime})\right|=\sup_{x^{\prime}\in\mathcal{X}}\sup_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}\left|f(x)-\widehat{f}(x^{\prime})\right|.$ (3.1)
@@
$L_{\Delta_{n}}^{*}(f)\triangleq\inf_{f^{\prime}}L_{\Delta_{n}}(f,f^{\prime})=\frac{1}{2}\sup_{x^{\prime}\in\mathcal{X}}\left[\sup_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)-\inf_{x\in(x^{\prime}-\Delta_{n})\cap\mathcal{X}}f(x)\right],$ (3.2)
@@
$f^{*}(x)=\frac{1}{2}\left[\sup_{x^{\prime}\in(x-\Delta_{n})\cap\mathcal{X}}f(x^{\prime})+\inf_{x^{\prime}\in(x-\Delta_{n})\cap\mathcal{X}}f(x^{\prime})\right],\quad x\in\mathcal{X}.$ (3.3)
@@ 3.2 Minimax rates of convergence
In this subsection, our aim is to establish the minimax rates of convergence for the sup-norm risk under the adversarial attacks. We propose an adversarial plug-in procedure to achieve the minimax optimal rates, which is derived from a minimax optimal estimator in the corresponding standard setting.
@@
Subsequently, we formulate an adversarial plug-in estimator:
$\widehat{f}_{\mathrm{PI}}(x)=\frac{1}{2}\left[\sup_{x^{\prime}\in(x-\Delta_{n})\cap\mathcal{X}}\widetilde{f}(x^{\prime})+\inf_{x^{\prime}\in(x-\Delta_{n})\cap\mathcal{X}}\widetilde{f}(x^{\prime})\right],\quad x\in\mathcal{X}.$ (3.4)
The performance of the adversarial plug-in estimator $\widehat{f}_{\mathrm{PI}}(x)$ clearly depends on the construction of $\widetilde{f}$. The following theorem first provides an upper bound for the adversarial risk of $\widehat{f}_{\mathrm{PI}}(x)$ considering a general $\widetilde{f}$. Additionally, Theorem 2 establishes minimax upper bounds when specific choices of $\widetilde{f}$ are adopted.
@@
$R_{\Delta_{n}}(f,\widehat{f}_{\mathrm{PI}})\leq R(f,\widetilde{f})+L_{\Delta_{n}}^{*}(f),$ (3.5)
@@
The following lower bound results show that the adversarial plug-in estimator based on $\widetilde{f}$ with (3.6) is in fact minimax rate optimal.
@@
In summary, Theorems 23 together establish the minimax rates of convergence for nonparametric regression under the adversarial attacks,
@@
Therefore, (3.10) addresses Question Q.1 raised in Section 2, showing that the adversarial minimax rate is jointly determined by the standard minimax rate and the largest ideal loss in $\mathcal{F}$. Regarding Question Q.2, we establish that if $\widetilde{f}$ is minimax optimal in the sense that $\sup_{f\in\mathcal{F}}R(f,\widetilde{f})\asymp\inf_{\widehat{f}}\sup_{f\in\mathcal{F}}R(f,\widehat{f})$ under the standard setting, then the adversarial plug-in estimator $\widehat{f}_{\mathrm{PI}}$ based on $\widetilde{f}$ is minimax optimal in terms of the adversarial risk. To the best our knowledge, (3.10) is the first minimax result in adversarial learning for the general regression setting. Our bounds are modular and can be applied to many models by computing the sup-norm convergence and the ideal adversarial loss in the target function class.
@@ 4 Applications
The marginal distribution $\mathbb{P}_{X}$ admits a density function that is lower bounded away from 0 and upper bounded by a positive constant on $\mathcal{X}$.
Assumption 1 ensures that the covariates $X$ are more or less evenly distributed over the compact support $[0,1]^{d}$. As a result, there are sufficiently many observations around any point in the support, allowing for the construction of well-behaved estimators for the regression function in the sup-norm loss. This assumption is standard in nonparametric regression with random design; see, for example, Condition 3’ in Stone, (1982) and Definition 2.2 in Audibert and Tsybakov, (2007). In addition, we further assume that the random error term is distributed according to a centered Gaussian distribution, which is the scenario where the known minimax theory in sup-norm can apply (see, e.g., Stone,, 1982; Bertin, 2004b, ; Gaïffas,, 2007).
@@ 4.1 Isotropic Hölder class
$\inf_{\widehat{f}}\sup_{f\in\mathcal{F}_{1}(\beta,L)}R_{\Delta_{n}}(f,\widehat{f})\asymp\left(\frac{\log n}{n}\right)^{\frac{\beta}{2\beta+d}}+C_{d}r_{n}^{1\wedge\beta},$ (4.3)
where $C_{d}\leq Cd^{k/2}$ is a constant not depending on $n$.
@@
The determination of the rate of $\sup_{f\in\mathcal{F}_{1}(\beta,L)}L_{\Delta_{n}}^{*}(f)$ is provided in Section S1 of the Supplementary Material.
@@
The quantity $r_{n}$ in (4.2) measures the length of the longest line segment contained in the set $\Delta_{n}$, and it may depend on the sample size $n$. The condition imposed on $\Delta_{n}$ is quite mild, which is satisfied by the $\ell_{p}$-ball: $B_{p}^{q_{n}}\triangleq\{\delta\in\mathbb{R}^{d}:\|\delta\|_{p}\leq q_{n}\}$, $0<p\leq\infty$, and the $\ell_{p}$-ball with the $\ell_{0}$-constraint: $B_{p}^{q_{n}}\cap\{\delta:\|\delta\|_{0}\leq s_{n}\}$. Note that there is an extensive body of prior work studying adversarial machine learning based on $\ell_{0}$ (Delgosha et al.,, 2024), $\ell_{2}$ (Bhattacharjee and Chaudhuri,, 2020; Bhattacharjee et al.,, 2021), and $\ell_{\infty}$ attacks (Athalye et al.,, 2018; Marzi et al.,, 2018). However, these analyses focus on the specific attacks and lack general applicability. In contrast, the result in Example 1 sheds theoretical insight on the adversarial robustness under the general $\ell_{p}$-attacks with $0<p\leq\infty$. Specifically, when $\Delta_{n}=B_{p}^{q_{n}}$, we have $r_{n}=q_{n}$, and thus the minimax adversarial risk is given by
@@
$\inf_{\widehat{f}}\sup_{f\in\mathcal{F}_{1}(\beta,L)}R_{B_{p}^{q_{n}}}(f,\widehat{f})\asymp\left(\frac{\log n}{n}\right)^{\frac{\beta}{2\beta+d}}+C_{d}{q_{n}}^{1\wedge\beta},$ (4.5)
@@
which can be reached by the adversarial plug-in estimator (3.4) with $\widetilde{f}$ constructed by a suitably designed local polynomial estimator (see, e.g., Stone, (1982), Gaïffas, (2007), and Tsybakov, (2008)).
@@
The equation (4.5) shows that when $\beta<1$ and $q_{n}\lesssim(\log n/n)^{1/(2\beta+d)}$, the minimax rate in the adversarial sup-norm remains unchanged to the standard minimax rate (4.4). However, as the magnitude of perturbation increases, e.g., $q_{n}\gtrsim(\log n/n)^{1/(2\beta+d)}$, the minimax risk has the order $q_{n}^{\beta}$. When $\beta\geq 1$ and the functions in $\mathcal{F}_{1}(\beta,L)$ become smoother, the critical radius $q_{n}$ for the phase transition is $(\log n/n)^{\beta/(2\beta+d)}$. It is also worth noting that the norm parameter $p$, which controls the shape of the perturbation set $B_{p}^{q_{n}}$, does not affect the adversarial minimax rates in this example. This is because the adversarial risk $R_{B_{p}^{q_{n}}}(f,\widehat{f})$ is defined via the worst-case perturbation within the ball $B_{p}^{q_{n}}$, and the minimax risk considers the worst-case adversarial risk over all functions in the function class. In fact, the maximum adversarial risk over the function class $\mathcal{F}_{1}(\beta,L)$ is attained at certain functions evaluated at points $x^{\prime},x$ satisfying $\|x^{\prime}-x\|_{2}=q_{n}$. Therefore, in the minimax sense, $p$ does not influence the adversarial minimax risk. However, in other regression function classes of interest, the shape of the perturbation may have an effect on the robustness of a given estimator; see Section 4.2 for further discussion.
Remark 2.
In this paper, we primarily focus on the adversarial sup-norm as the robustness performance measure. Using the uniformity of the sup-norm loss, we can derive the following upper bound on the adversarial $L_{2}$-loss
$\bar{L}_{\Delta_{n}}(f,\widehat{f})\triangleq\int_{\mathcal{X}}\sup_{\begin{subarray}{c}\delta\in\Delta_{n}\\ x+\delta\in\mathcal{X}\end{subarray}}\left|f(x)-\widehat{f}(x+\delta)\right|^{2}\mathbb{P}_{X}dx\lesssim L_{\Delta_{n}}^{2}(f,\widehat{f}),$
under the assumption that $\mathcal{X}$ is a compact set and $\mathbb{P}_{X}$ satisfies Assumption 1. Based on this relation and (4.3), we can also derive an upper bound on the minimax adversarial risk under $L_{2}$-loss over the isotropic Hölder class:
$\left(\frac{\log n}{n}\right)^{\frac{2\beta}{2\beta+d}}+C_{d}{r_{n}}^{2(1\wedge\beta)}.$
It remains to be seen if this is the minimax optimal rate.